# There is no universal separable Banach algebra

### v1

@@ 2. Preliminaries
First, we present a single proof that establishes both assertions (A) and (B) simultaneously. Let $B$ be an arbitrary separable commutative Banach algebra, and write $R:=\operatorname{rad}(B)$ for its Jacobson radical. Since $B$ is separable, so too is $R$. Set
$G:=R\widehat{\otimes}_{\pi,\mathrm{s}}R,$
the symmetric projective tensor product of $R$ with itself. Then $G$ is a separable Banach space.
By the Johnson–Szankowski theorem [4, Theorems 2.5–2.6], there exist separable Banach spaces $U$ and $Y$ and a compact operator $S:U\to Y$ which does not factor through the separable space GG. [^1] By a consequence of Johnson–Szankowski [4, Thm. 2.5] (see also [3] for background), for every separable Banach space $G$ there exist separable Banach spaces $U$ and $Y$ and a compact operator $S:U\to Y$ such that $S$ does not belong to the operator ideal of maps factoring through $G$ (i.e., $S$ does not factor through $G$).
@@
We now construct a separable commutative test algebra that cannot embed into $B$. Write $X:=U\oplus_{1}\mathbb{C}$ (the $\ell_{1}$-direct sum), and define a symmetric bilinear map $\beta_{U}:X\times X\to U$ by
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda u^{\prime}+\lambda^{\prime}u$
@@
for all $(u,\lambda),(u^{\prime},\lambda^{\prime})\in X$. Let $E:=X\widehat{\otimes}_{\pi,\mathrm{s}}X$, and let $\widetilde{\beta}^{\,\mathrm{sym}}_{U}:E\to U$ be the unique bounded linearisation of $\beta_{U}$. By Lemma 2.13, the map $\widetilde{\beta}_{U}$ is surjective with bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$, and the compact operator
$T_{0}:=S\circ\widetilde{\beta}^{\,\mathrm{sym}}_{U}:E\longrightarrow Y$
does not factor through $G$ (for if it did, then $S=T_{0}\circ\sigma$ would factor through $G$, contradicting our choice of $S$). Now define a symmetric bilinear map $\beta:X\times X\to Y$ by composition:
$\beta:=S\circ\beta_{U}.$
Then the linearisation of $\beta$ is precisely $\widetilde{\beta}^{\,\mathrm{sym}}=S\circ\widetilde{\beta}^{\,\mathrm{sym}}_{U}=T_{0}$. By Proposition 2.9, the $\ell_{1}$-direct sum
$A:=A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y,$
equipped with the multiplication
$(\alpha,x,y)\cdot(\alpha^{\prime},x^{\prime},y^{\prime}):=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr),$
is a commutative, unital Banach algebra (possibly after an equivalent renorming) with radical $\operatorname{rad}(A)=X\oplus Y$. Since $U$, $Y$, and hence $X$ are all separable, the algebra $A$ is separable.
Suppose, towards a contradiction, that there exists an embedding $\phi:A\hookrightarrow B$. (In case (A), we require $\phi$ to be contractive; in case (B), we merely require $\phi$ to be bounded. In either case, $\phi$ is an injective homomorphism that is a topological isomorphism onto its closed range in $B$.) Write
$i:=\phi|_{X}:X\longrightarrow R\qquad\text{and}\qquad j:=\phi|_{Y}:Y\longrightarrow R$
for the restrictions of $\phi$ to $X$ and $Y$, respectively. By Lemma 2.10, we have the factorisation identity
(2.3) $j\circ\widetilde{\beta}^{\,\mathrm{sym}}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i)\quad:\quad E=X\widehat{\otimes}_{\pi,\mathrm{s}}X\longrightarrow R,$
where $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\widehat{\otimes}_{\pi,\mathrm{s}}R\to R$ is the symmetric linearisation of multiplication in $R$. Moreover, since $\phi$ is an embedding, Lemma 2.10 also tells us that $j:Y\to R$ is bounded below (and in particular injective). Thus $j$ is a topological isomorphism of $Y$ onto the closed subspace $\operatorname{ran}(j)\subseteq R$, and the inverse map $(j|_{\operatorname{ran}j})^{-1}:\operatorname{ran}(j)\to Y$ is bounded.
Using (2.3) and $\widetilde{\beta}^{\,\mathrm{sym}}=T_{0}$ we have
$j\circ T_{0}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i)\ :\ E\longrightarrow R,$
@@
so $j\circ T_{0}$ factors through $G=R\widehat{\otimes}_{\pi,s}R$ (via $E\xrightarrow{i\otimes_{\mathrm{sym}}i}G\xrightarrow{\ \widetilde{\mu}_{R}^{\,\mathrm{sym}}\ }R$). By Lemma 2.13 we have a bounded right inverse $\sigma:U\to E$ of $\widetilde{\beta}^{\,\mathrm{sym}}_{U}$ with $S=T_{0}\circ\sigma$, hence
@@
also factors through $G$. Since $\phi$ is an embedding, $j:Y\to R$ is injective and bounded below, so the inverse $(j|_{\operatorname{ran}j})^{-1}:\operatorname{ran}j\to Y$ is bounded. Regard $j\circ S$ as an operator with codomain $\operatorname{ran}j$ (a Banach space with the inherited norm). Post-composing by $(j|_{\operatorname{ran}j})^{-1}$ yields
$S\;=\;(j|_{\operatorname{ran}j})^{-1}\circ(j\circ S),$
and factorisations are stable under bounded post-composition. Therefore $S$ also factors through $G$, contradicting the choice of $S$. This completes the proof of (A) and (B).
The argument for parts (C) and (D) follows the same overall strategy but with two crucial differences: we work with the unsymmetrised projective tensor product $B\widehat{\otimes}_{\pi}B$, and we avoid any reliance on radical structure. The test algebra construction uses a non-symmetric bilinear form, and the factorisation proceeds directly through multiplication in $B$.
Fix a separable Banach algebra $B$ and set $G:=B\widehat{\otimes}_{\pi}B$, which is separable. Again, by the Johnson–Szankowski theorem, there exist separable spaces $U,Y$ and a compact operator $S:U\to Y$ which does not factor through $G$. Form $X:=U\oplus_{1}\mathbb{C}$ and $\beta_{U}$ as in Lemma 2.14; let $E:=X\widehat{\otimes}_{\pi}X$ and $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, which does not factor through $G$. Define $\beta:=S\circ\beta_{U}:X\times X\to Y$ and let $A:=A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y$, a separable Banach algebra by Proposition 2.9. Assuming $B$ is universal, there is an embedding $\phi:A\hookrightarrow B$. Write $i:=\phi|_{X}$ and $j:=\phi|_{Y}$. By Lemma 2.11,
$j\circ T_{0}\;=\;\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i).$
Set $P:=i\otimes_{\pi}i:E\to G$ and $Q:=\widetilde{\mu}_{B}:G\to B$; then
$j\circ T_{0}=Q\circ P,$
so $j\circ T_{0}$ factors through $G=B\widehat{\otimes}_{\pi}B$. By Lemma 2.14 we have $j\circ S=(j\circ T_{0})\circ\sigma=Q\circ(P\circ\sigma)$, hence $j\circ S$ also factors through $G$. Since $\phi$ is an embedding, $j$ is injective and bounded below; viewing $j\circ S$ with codomain $\operatorname{ran}j$ and post-composing with $(j|_{\operatorname{ran}j})^{-1}$ shows that $S$ factors through $G$, a contradiction. ∎

### v2

@@ 3. Proof of the main theorem
We present a unified proof that covers all four cases, (A)-(D). The argument proceeds by contradiction.
Let $B$ be an arbitrary separable Banach algebra, which we suppose to be a universal object in one of the clauses described. If we are addressing the commutative categories (A) or (B), we assume $B$ is commutative; otherwise, no such assumption is made. In all cases, set $G:=B\widehat{\otimes}_{\!\pi}B$, which is a separable Banach space as $B$ is. By the Johnson–Szankowski theorem [3], there exist separable Banach spaces $U,Y$ and a compact operator $S:U\to Y$ which does not factor through $G$.
@@
Our strategy is to construct a single separable test algebra $A$ that cannot be embedded into $B$, regardless of whether $B$ is commutative. To this end, form $X:=U\oplus_{1}\mathbb{C}$ and define the symmetric bilinear map
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda u^{\prime}+\lambda^{\prime}u.$
@@
Let $E:=X\widehat{\otimes}_{\!\pi}X$ and let $\widetilde{\beta}_{U}:E\to U$ be its (unsymmetric) linearisation. By Lemma 2.7, $\widetilde{\beta}_{U}$ is surjective with a bounded right inverse $\sigma:U\to E$, and the composite operator $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$ does not factor through $G$.
Now, define the bilinear map $\beta:=S\circ\beta_{U}:X\times X\to Y$. Since $\beta_{U}$ is symmetric, so is $\beta$. We then construct the test algebra $A:=A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y$ using the multiplication from Proposition 2.5. As $\beta$ is symmetric, $A$ is a commutative separable Banach algebra.
Suppose, towards a contradiction, that there exists a bounded homomorphic embedding $\phi:A\hookrightarrow B$. (Note that if a contractive embedding exists, it is also a bounded one, so this assumption covers all cases.) Write $i:=\phi|_{X}$ and $j:=\phi|_{Y}$. By Lemma 2.6, the fact that $\phi$ is a homomorphism forces the following identity on the linearisations:
$j\circ\widetilde{\beta}\;=\;\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i)\quad:\quad E\longrightarrow B.$
@@
Recalling that $\widetilde{\beta}=S\circ\widetilde{\beta}_{U}=T_{0}$, this means $j\circ T_{0}$ factors through $G$. Since $\widetilde{\beta}_{U}$ has a bounded right inverse $\sigma$, we can compose with it to find that
@@
also factors through $G$. This factorisation is illustrated below.
[…]
As $\phi$ is an embedding, the map $j$ is a topological isomorphism onto its range, so it has a bounded inverse $(j|_{\operatorname{ran}j})^{-1}:\operatorname{ran}j\to Y$. Pre-composing the factorisation of $j\circ S$ with this inverse shows that $S$ itself must factor through $G$. This contradicts our choice of $S$.
The contradiction implies that no such embedding $\phi:A\to B$ can exist.
  • In cases (A) and (B), we assumed $B$ was a universal separable commutative Banach algebra. We have constructed a separable commutative Banach algebra $A$ that does not embed into $B$, a contradiction.
  • In cases (C) and (D), we assumed $B$ was a universal separable Banach algebra for all separable Banach algebras. We have constructed a specific separable (and, as it happens, commutative) Banach algebra $A$ that does not embed into $B$, which again is a contradiction.
Since our choice of $B$ was arbitrary, no such universal object exists in any of the specified categories. The argument relied only on the boundedness of the embedding, so it holds equally for contractive embeddings. This completes the proof of (A)–(D). ∎